Write an equation:
1. The sum of half of a number and 8 less than the number is 25.
n + (n-8) = 25 (?)
2. A triangle has two sides equal in length and a third side 5 in. longer than half the length of each of the other two sides. If the perimeter of the triangle is 50 in., how long is each side?
2x + (x + 5) = 50 (?)

Answer by stanbon(75887)

1. The sum of half of a number and 8 less than the number is 25.
Let the number be "x":
Equation:
(1/2)x + (x-8) = 25
Multiply thru by 2 to get rid of the fraction:
x + 2x-16 = 50
3x = 66
x = 22
-----------------------------------

2. A triangle has two sides equal in length and a third side 5 in. longer than half the length of each of the other two sides. If the perimeter of the triangle is 50 in., how long is each side?
----
Equation:
2x + ((x/2 + 5) = 50
2x + x/2 + 5 = 50
(5/2)x = 45
x = (2/5)45
x = 18 (length of each of the equal sides)
(x/2) + 5 = 9+5 = 14 (length of the 3rd side.
===============================================
Cheers,
Stan H.